I don't understand this, can someone help?
Write the equation of the line passing through each of the given pairs of points. Write your result in slope-intercept form, where possible.
(-4, 1) and (3, 1)

Answer by stanbon(75887)

Write the equation of the line passing through each of the given pairs of points. Write your result in slope-intercept form, where possible.
(-4, 1) and (3, 1)
---------------------
slope = (1-1)/(3--4) = 0/7 = 0
---
Form: y = mx + b
Substite for y,x, and m ; solve for "b":
1 = 0*3 + b
b = 1
----
Equation:
y = 0x + 1
y = 1
=========================
Notice that both of your points have y = 1.
So the equation is a horizontal line where y = 1
=========================
Cheers,
Stan H.